OpenStudy (anonymous):

x^2-ax-1=0, x^3+1/x^2+x^2-1/x^3+1 what is a?

OpenStudy (anonymous):

I asked wolfram but still i don't get its answer.

OpenStudy (jack1):

x^3+1/x^2+x^2-1/x^3+1 =??? you have 2 unknowns so dont you need 2 equations to solve it?

OpenStudy (jack1):

ooo... shiny, cheers

OpenStudy (anonymous):

wait sorry i was trying to say put this equation in terms of a and do factoring

OpenStudy (jack1):

ah, k, in that case: x^2-ax-1=0 -1 = x^2-ax a= x- (1/x)

OpenStudy (anonymous):

hmmm. What do you mean by k?

OpenStudy (jack1):

so... shortened version of okay, not a constant

OpenStudy (jack1):

sorry

OpenStudy (anonymous):

ok

OpenStudy (jack1):

(x^3+1)/(x^2+x^2-1)/(x^3+1) is this eqn 2? x^3+(1/x^2)+x^2-(1/x^3)+1 ...or this?

OpenStudy (anonymous):

so should i make it like (x-1)(x^2+1+1/x^2)(x^+1+1/x^1)

OpenStudy (jack1):

x^3+(1/x^2)+x^2-(1/x^3)+1 =x^3-(1/x^3)+x^2+(1/x^2)+1 so a^2= x^2+(1/x^2) +2 therefore take that out of above leaves remainder: x^3-(1/x^3)-1

OpenStudy (jack1):

so now u have: a^2 + x^3-(1/x^3)-1 so just keep going untill you get =a^+ 3a +a^3 and remainder of +3

OpenStudy (jack1):

so reduces to =a^2+ 3a +a^3+3

OpenStudy (anonymous):

oh thanks.